a school administrator claims that less than 50% of the students of the school are dissatisfied by the community cafeteria service. Test this claim by using sample data obtained from a survey of 500 students of the school where 54% indicated their dissatisfaction of the community cafeteria service. use a 0.05.


1
Expert's answer
2021-06-09T16:35:12-0400

Answer:-

P: Population proportion dissatisfied by the community cafeteria service .

HO: P =0.5

H1=P<0.5

"\hat{P}:" Sample proportion dissatisfied by community cafeteria service

"\hat{P}=0.54"

"Z=\frac{\hat{P}-P}{\sqrt{\hat{P}(1-\hat{P})\over n}}" ~ N(0,1)


"Z=\frac{0.54-0.5}{\sqrt{(0.54)(0.46)}\over 500}" "\implies" "\frac{0.04}{0.022}"

"Z=1.82"

Level of significance "\alpha=0.05"

"Z_{\alpha}" (onetail)=1.64

"Z>Z_{\alpha}" at 5% level of significance , thus Ho will be rejected.

Hence school administer claim of "(P<0.5)" will be accepted.